OpenStudy (anonymous):

Solve the equation (x+1)(x+2)(x+3)(x+4)=360

OpenStudy (anonymous):

just tell me the starting i'll do the rest

OpenStudy (anonymous):

is anyone helpin ??

OpenStudy (anonymous):

For this equation, you can expand that left side then move the 360 over to the left. Then you can use synthetic division to try and factor to find the solution.

OpenStudy (anonymous):

(y+4)(y+6)=360

OpenStudy (anonymous):

where did the x's go? expand: (x+1)(x+2)(x+3)(x+4)

OpenStudy (anonymous):

I used a trick

OpenStudy (anonymous):

oh... sorry... i'm not familiar with that trick.... can you show me how?

OpenStudy (anonymous):

rearranged the terms (x+1)(x+4)(x+3)(x+2)=360 (x^2+5x+4)(x^2+5x+6)-360=0 let y=x^2+5x (y+4)(y+6)-360=0

OpenStudy (anonymous):

ok... you'll still need to expand that left side involving y's though....

OpenStudy (anonymous):

I solved the equation and got \[{2,-7,\frac{ -5\pm \sqrt{-71} }{ 2 }}\]

OpenStudy (anonymous):

oh... you're also looking for complex roots? i was only looking for real ones.. and those first two are the ones i got.... nice work!!!

OpenStudy (anonymous):

thanks